How we can calculate pressure at different temp, e.g butane at 20 C & it's vapour pressure (absolute pressure) 220 kPa similarly, at 40 C is 360 kPa, at 45C is 385 kPa & 55C is 580 kPa. I want to calculate more pressure values at different temp.
1
Expert's answer
2016-01-10T06:58:30-0500
According to Gay-Lussac’s law the pressure of gas is proportional to the absolute temperature upon constant volume.
This can be expressed by the equation:
P=K×T, where P – the pressure, K – constant, T – the absolute temperature [T = 273 + t(in Co)].
If we need to calculate more pressure values at different temperature, the K parameter should be found:
220 kPa = K(273+20), K = 0.75085,
360 kPa = K(273+40), K = 1.1501,
385 kPa = K(273+45), K = 1.2106,
580 kPa = K(273+55), K = 1.7683
As we can see the K values are different. Calculating its average value we get an equation, which gives the pressure at given temperature:
Kav = 1.22

Answer: Thus, P = 1.22×T.